If x^2 - 5x + 6 = (x - h)^2 + k, what is the value of k?
(A) - 25/4 (B) -5/2 (C) -1/4 (D) 0 (E) 6

Answer by stanbon(75887)

If x^2 - 5x + 6 = (x - h)^2 + k, what is the value of k?
----
x^2-5x+6 = x^2-2hx+h^2+k
---
Equate the coefficients:
-2h = -5
h^2+k = 6
------
h = 5/2
Substitute and solve for "k":
(5/2)^2+k = 6
k = 6-(25/4)
---
k = -1/4
==============
Cheers,
Stan H.
======================

(A) - 25/4 (B) -5/2 (C) -1/4 (D) 0 (E) 6